OpenStudy (anonymous):

statistics question: john is going on a 5-day holiday to Costa Packet. the travel brochure says that on average 5 out of every 7 days are sunny at costa packet. each days weather is independent of the weather on all preceding days. a)there are two values, n and p, that you need to use in binomial distribution, write down the value of n and the value of p. b) calculate the probability that john has 2, or less, sunny days on his holiday? you may use: (p+q)⁵ = p⁵ + 5p⁴q + 10p³q² + 10p²q³ + 5pq⁴ + q⁵ c)what is the most likely number of sunny days that john will get on his holiday?

OpenStudy (anonymous):

well in the front of your textbook you should have a z-score chart and that is exactly what this is is a z-score question first of all z-score counts how many standard deviations a number is away from the mean.. so what you need to do is take the mean-(number given) / standard dev this will give you the z-score and from there you look that number up on the chart and it will tell you the percentage (Going down..so if it says higher/more/greater than you need to subtract the number from 100).. firstly a) less than 600$ z-score = 680-600 / 138 = -.5797 standard devs away from the mean.. that zscore gives you a percentage of 30.5% b)more than 700$ z-score= 680-700/138 = .1449 standard devs away from the mean.. that zscore gives you a percentage of 92.65% c)between 600$ and 700$ for these you take the lower zscore percentage and subtract it from the higher zscore percentage. 600$ = 30.5% 700$ = 92.65% 92.65%-30.5% = 62.15% those are your three answers hope i cleared things up a little :) you can always use an online zscore chart. Source(s): http://www.epatric.com/documentation/sta

OpenStudy (anonymous):

but i got n=7 and p=5/7 :/

OpenStudy (anonymous):

u r correct

OpenStudy (anonymous):

me?

OpenStudy (anonymous):

ya!

OpenStudy (anonymous):

would youu be able to help me with the rest please? :)

OpenStudy (anonymous):

once youve answered it

OpenStudy (anonymous):

what on earth is going on here?

OpenStudy (anonymous):

a lot of words again that have nothing to do with the question. this character needs to be stopped

OpenStudy (anonymous):

first off you are right about p since it says sunny five out of seven days \(p=\frac{5}{7}\) but you are mistaken about n n is the number of days on the trip (the number of elements in the "sample space") and it is 5

OpenStudy (anonymous):

@satellite73 , couldn't agree with you more !

OpenStudy (anonymous):

so we have \[n=5,p=\frac{5}{7}, 1-p=\frac{2}{7}\]

OpenStudy (anonymous):

it is one thing to give a wrong answer because you are mistaken i have certainly done that. it is quite another to copy and paste some non related garbage and post it as an answer, then say something random

OpenStudy (anonymous):

so now will it follow binomial distribution?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@islacameron if you are still there, we can complete this problem if you like. once we have \[n=5,p=\frac{5}{7}, q=1-p=\frac{2}{7}\]

OpenStudy (anonymous):

two or less sunny days means \[P(x=0)+P(X=1)+P(X+2)\]

OpenStudy (anonymous):

you have \[(\frac{2}{7})^5+6(\frac{2}{7})^4\times \frac{5}{7}+10(\frac{2}{7})^3\times (\frac{5}{7})^2\]

OpenStudy (anonymous):

normally you would have to compute the binomial coefficents but they were given to you in this line above \[(p+q)⁵ = p⁵ + 5p⁴q + 10p³q² + 10p²q³ + 5pq⁴ + q⁵ \]

OpenStudy (anonymous):

Easiest of all c)what is the most likely number of sunny days that john will get on his holiday? Answer -->5

OpenStudy (anonymous):

@islacameron if you have any questions let me know, i will be happy to answer about this problem. ignore @Rohangrr posts, they are almost always random

OpenStudy (anonymous):

i htink most likely sunny days needs to be computed. sunny all five days has probability \((\frac{5}{7})^5=.186\) rounded, whereas 4 sunny days is \(5\times (\frac{5}{7})^4\times \frac{2}{7}=.372\) rounded

OpenStudy (anonymous):

evidently the probability of 4 is larger than the probability of all five we can also compute the probability of exactly 3 sunny days via \(10\times (\frac{5}{7})^3(\frac{2}{7})^2=.297\) rounded

OpenStudy (anonymous):

unless you were talking about someone that gave the answer then in xthat case im am terribly sorry and take all of that back! :)

OpenStudy (anonymous):

if you are addressing me, my comments about random answers were directed to rohangrr's post, not your question i hope i answered your question completely, and if you have any question about what i wrote i would be happy to explain in detail

OpenStudy (anonymous):

ahh well yeah in that case i apologise and that you a lot for your help and answer, sorry

OpenStudy (anonymous):

that nonsense starting with "z scores" has nothing to do with the problem

OpenStudy (anonymous):

did you get that \(n=5, p=\frac{5}{7}\) ?

OpenStudy (anonymous):

yeahh, thanks

OpenStudy (anonymous):

and the rest as well for the binomial distribution?

OpenStudy (anonymous):

yeah thanks :)

OpenStudy (anonymous):

ok, good.